# Autonomous Issue Resolver: Towards Zero-Touch Code Maintenance

### v2

@@ 6.1 Agent environment
Figure 6.2: The AIR Multi-Agent Architecture. The system employs a decoupled ”Plan-Navigate-Execute” loop to manage context window limitations. The Context Agent (pink) operates exclusively within the graph domain to localize faults without reading file content. Once the fault subgraph is identified, the Maintenance Agent (yellow) formulates a high-level repair plan, which is translated into concrete syntax edits by the Editor Agent (gray). This separation prevents the ”Semantic Trap” by ensuring the planner is conditioned only on causally relevant data lineages rather than raw repository noise.
We model agents whose internal state is a structured view over the DTG.
@@
5 graph_view: SubGraph # local neighborhood around current_node_id
@@
The agent is equipped with a small set of graph-oriented tools (Table 6.2). These tools abstract away file-system operations and present the agent with a clean graph navigation API.
@@
Table 6.2: Available Navigation and Testing Functions
@@ 7.1 Benchmarks and Metrics
To test our tool we use the following benchmarks.
SWE-bench family. To test the data-first hypothesis, we outline an evaluation protocol focusing on repair effectiveness, localization efficiency, and context utilization. We evaluate on the SWE-bench suite (Full, Verified, Lite; using metric: %Resolved with repository tests[32, 15].
APR. QuixBugs (Java/Python) [33] and HumanEvalFix.
Industrial tickets. A stratified sample of $N$ internal issues across the services.
Primary metrics: %Resolved (tests pass), MTTR (hours), regression rate (%), reviewer effort (minutes/patch), CI cost (compute \$), and change risk score (static+dynamic).
@@ 7.2 Baselines
We employ the standard baselines, popular in coding tasks: SWE-agent [11], OpenHands [12], AutoCodeRover [13], and an ablated AIR without RL (AIR w/o RL) and without DTG. The numbers for the baselines are taken from the original papers and the benchmark website. [32]
@@ 7.3 Compute Budget and Settings
Uniform wall-clock budget per issue (e.g., 60–120 minutes), identical tool access, and fixed seeds. We report median over three runs and 95% CIs.
7.4 Close-to-life experiment
On the top of classic benchmarks, we simulated a realistic scenario, where the AI agent is tasked with resolving reported software issues and its output is reviewed by a professional engineer. We selected 10 open-source Python repositories to represent a spectrum of complexity, from low-dependency utilities (jsonpickle) to high-complexity image processing libraries involving C-extensions (Pillow). The selection criteria required repositories to have active maintenance, >1k>1k stars, and a history of well-documented issue resolution to serve as the ground truth. ”Complexity” is qualitatively assessed based on dependency depth and code coupling (see Table 7.3). The evaluation utilized a curated set of $N=100$ coding tasks, split into 50 easy, 40 medium and 10 high-complexity challenges.
Table 7.3: Github repositories used in our tests. Total issues column indicates the number of reported issues in repository, snapshot of 12/05/2025
@@ 7.4 Close-to-life experiment
Baseline values are from the cited papers/sites; update to latest leaderboard snapshots.
@@
    Reported %Resolved on SWE-bench at time of writing [32] (see the data from the original paper in [11]).
@@
    For current SWE-bench leader-boards and variants (Full/Verified/Lite), see the website [32, 15].
@@ 8.1 Components
We ablate (i) RL control (AIR w/o RL), and (ii) DTG (AIR w/o Graph).
@@ 9.1 Ablations: What matters and why
Table 8.5 shows that removing RL control (AIR w/o RL) reduces %Resolved from 87.1 to 82.5 while shrinking MTTR. In this case, the shorter MTTR reflects premature termination rather than efficient resolution: an agent gives up quicker without policy guidance.
Removing DTG (w/o Graph) yields a mid-range drop (%Resolved 56.2), indicating repository-scale context is a major contributor.
Overall, RL and Graph are the largest drivers of resolution rate.
@@ 9.2 Comparison with commercial tools
Based on the data provided, here are the calculated statistics for each model’s performance across the 10 projects (100 total issues). This test is less reliable than SWE-bench, because of several factors. First, we have a much lower number of datapoints. In particular, we pre-selected 10 issues (out of 100s for some repos) on the basis of their resolvability - we wanted to shortlist the purely coding issues that can be solved by just applying best coding practice and debugging skills. That introduces some biases.
The context for each case was built artificially, including all files that were referenced in final PR that solved the issue and which played the role of ground truth.
For evaluation we used a consistent and objective evaluation methodology. In particular, we awarded score +1 for solving the issue (all requirements in the ticket are satisfied) and all unit tests are passed. We did not penalize for time, cost or failed solutions. The Figure 9.3 depict the relative performance of each model, all data is aggregated in Table 9.6, with summary statistics in Table 9.7
@@
Figure 9.3: Testing results on GitHub Issues. Numbers indicate how many issues were resolved. We randomly selected 10 issues from each repository.
Table 9.6: Issue resolution on five OSS GitHub repos (10 issues each; same wall-clock + tool budget). Numbers indicate how many issues were resolved.
@@
Table 9.7: Performance statistics per model across 10 projects (100 total issues).
@@
  • AIR (this work) significantly outperformed the other models, solving 70% of all issues. It was the only model to solve at least 4 issues in every single project.
@@ 9.2.1 Performance Analysis and the ”Semantic Gap”
The results in Figure 9.3 and Table 9.6 reveal a strong correlation between the explicitness of data flow and the agent’s success rate. AIR achieves a 70% resolution rate overall, showing particular strength in ”loose-coupling” architectures like requests and urllib3 (8/10 and 9/10 respectively). In these repositories, logic errors are often distributed across multiple abstraction layers - a scenario where the DTG’s ability to trace data lineage allows the agent to ”hop” over irrelevant middleware, a capability lacking in the file-based baselines.
@@ 9.2.2 Failure Case Study: The Pillow Repository
A notable deviation is observed in the Pillow repository, where commercial models scored 0/10 and AIR achieved only 4/10. We hypothesize this performance drop stems from the ”Language Boundary” problem. Pillow relies heavily on C-extensions for image processing. The current DTG implementation parses Python ASTs but treats C-bindings as opaque edges. Consequently, when data flows into a compiled extension, the lineage is broken, blinding the agent. This limitation highlights the necessity for future work in cross-language graph construction to support hybrid-runtime environments.
10 Conclusion and Future Roadmap
We have argued that the dominant file-first, control-centric view of repositories is misaligned with both human debugging practice and the causal structure of many real bugs. By re-centering our representation around data states and transformations, the Data-First Transformation Graph (DTG) offers a compact and causally meaningful substrate for LLM-based repair agents.
Our empirical results on SWE-bench Verified, where AIR achieved an 87.1% resolution rate, validate that shifting from text-based retrieval to graph-based navigation substantially reduces the search complexity of logic repair. However, our analysis of failure cases, particularly within the Pillow repository, highlights critical avenues for future development. We outline three primary directions for the evolution of AIR:
1. Breaking the Language Boundary (Polyglot DTGs): Our experiments revealed that the current DTG construction, while effective for pure Python, struggles with hybrid runtimes (e.g., Python wrappers around C/C++ extensions). Future iterations must implement cross-language graph builders that can trace data lineage through Foreign Function Interfaces (FFI). This involves unifying AST parsers across languages into a shared intermediate representation, allowing the agent to ”tunnel” through compiled extensions rather than treating them as opaque edges.
2. Dynamic Graph Refinement via RL: Currently, the DTG is static, constructed prior to the agent’s traversal. We propose a dynamic architecture where the agent can modifying the graph structure at runtime—collapsing well-tested subgraphs to save context or expanding ambiguous nodes for deeper inspection. By integrating this into the Reinforcement Learning policy, the agent can learn to optimize its own environment representation, effectively performing ”attention management” over the repository topology.
@@ 10 Conclusion and Future Roadmap
Transforming the LLM’s world model from a ”pile of files” to a navigable network of data flows, we move closer to truly autonomous systems capable of maintaining the invisible infrastructure of our digital civilization, precisely tuned to the inner logic of software. AIR represents the first step in this data-first transformation, bridging the gap between local code completion and systemic software reasoning.
Acknowledgments
The authors would like to thank Modal and lightning.ai that provided compute resources and free cloud credits for our experiments.
Appendix A Building the DTG from source files
@@ 27 ConsoleMessage(
Listing 1: A code snippet from poetry project (poetry/src/poetry/installation/chooser.py:125:197). Some non-essential lines were cut for brevity
The graph builder can extract the following raw constructs from this snippet of code (Table A.1), and eventually build a subgraph depicted in Figure A.1. We used a standard ast library for parsing python files, neo4j for storing nodes and edges, and ChromaDB as a vector storage tier for semantic search of relevant nodes/edges.
@@
Table A.1: Summary of extracted code constructs. We omit verbose metadata for breivity, such as file name, line number, type of variables, etc. All function arguments, function value and local variables will become the Data Nodes. Function name will become the basis for the edge. All formed structures will be stored in neo4j database, and the embeddings for underlying code will be stored in vector database (we use ChromaDB)
@@
Figure A.1: The reconstructed sample of Data Transformation Graph given the raw data to build Nodes and Edges from Table A.1.
Appendix B AI-assisted software tools
Table B.2 presents a comparison of market tier, interaction modality, and primary capabilities of today’s AI coding tools. While ”Professional” tier tools (e.g. Copilot, Cursor) dominate the interactive space via IDE integration, they operate primarily on a ”Human-in-the-Loop” synchronous model and hardly can be considered a serious AI technology. ”Specialized” agents (AlphaCode) excel at algorithmic synthesis but lack repository awareness. AIR is distinct in its focus on asynchronous, autonomous maintenance, targeting the ”Zero-Touch” quadrant.
Table B.2: The landscape of AI-assisted software tools (Q4’25). We distinguish market tier, modality, and granularity to expose differences that matter for maintenance automation.
Product Tier Modality Primary Capability Notes / Differentiators (incl. SOTA area)
@@ Appendix B AI-assisted software tools
GitHub Copilot [6] Professional IDE/Editor plugin Predictive completion, chat, refactors Strong ecosystem integration (PR comments, inline hints); SOTA for ergonomic code completion at scale.
@@
Amazon CodeWhisperer [7] Professional IDE plugin Completion + security hints Security-aware suggestions; policy scanning baked into cloud workflows.
@@
Tabnine [34] Professional IDE plugin Local+cloud completion Emphasis on on-device privacy modes; enterprise control.
@@
Cursor [8] Professional AI-first IDE Conversational edit, repo search Editor tuned around agentic edits; tight repo context UX.
@@
Claude Code [17] Professional Chat+tools Conversational code reasoning Long-context code reasoning; strong natural language guidance.
@@
OpenAI Codex (legacy) Professional API/chat (retired) Codegen foundation model Historic baseline; catalyzed tool ecosystem.
Sourcegraph [35] Professional IDE plugin Autonomous reasoning, comprehensive code editing Advanced code assistant with many autonomous features
B44 (builder) [36] Basic Web builder App scaffolding from prompts Fast MVP creation; limited depth, human follow-through required.
@@
Lovable [37] Basic Web builder No/low-code + AI Rapid prototyping; non-expert friendly; limited repo-scale change mgmt.
Replit (Ghostwriter) [38] Basic Web IDE In-editor assist Lightweight, education/startup focus; good DX, modest repo semantics.
AlphaCode family [28] Specialized Research system Competition-style codegen SOTA in competitive programming; not integrated with CI/CD or repos.
@@
AlphaDev [27] Specialized Research system Algorithm discovery/optim. SOTA algorithm discovery; narrow domain, non-general maintenance.
@@
Diffblue Cover Specialized JVM toolchain Unit test synthesis Industrial unit test generation; complements but does not fix issues.
@@
Getafix/Infer lineage [26] Specialized Internal/OSS tools Learned fix patterns + static analysis Pattern mining for APR; depends on rich bug corpora; limited autonomy.
@@
Factory.ai [39] Specialized Cloud integration Agent-native Software development Corporate level maintenance tool

### v3

@@ 5.1 Agent environment
Figure 2: The AIR Multi-Agent Architecture. The system employs a decoupled ”Plan-Navigate-Execute” loop to manage context window limitations. The Context Agent (pink) operates exclusively within the graph domain to localize faults without reading file content. Once the fault subgraph is identified, the Maintenance Agent (yellow) formulates a high-level repair plan, which is translated into concrete syntax edits by the Editor Agent (gray). This separation prevents the ”Semantic Trap” by ensuring the planner is conditioned only on causally relevant data lineages rather than raw repository noise.
Agents are modeled with an internal state represented as a structured view over the DTG.
@@
5 graph_view: SubGraph # local neighborhood
@@
The agent is equipped with a small set of graph-oriented tools (Table 2). These tools abstract away file-system operations and present the agent with a clean graph navigation API.
@@
Table 2: Available Navigation and Testing Functions
@@ 6.1 Benchmarks and Metrics
The evaluation uses two public benchmark families and one exploratory repository-level study.
SWE-bench. The system is evaluated on SWE-bench Verified and SWE-bench Lite using %Resolved with repository tests as the primary metric [17, 14].
APR. HumanEvalFix is included as an additional APR reference benchmark in Table 4.
OSS study. An exploratory repository-level study is conducted on 10 open-source Python repositories, with 10 issues sampled per repository.
Primary metrics. The main reported metric is %Resolved for benchmark tasks, together with issue-resolution counts and summary statistics for the OSS study.
@@ 6.2 Baselines
The comparison includes the standard baselines, popular in coding tasks: SWE-agent [21], OpenHands [19], AutoCodeRover [25], and an ablated AIR without RL (AIR w/o RL) and without DTG. The numbers for the baselines are taken from the original papers and the benchmark website. [17]
@@ 6.3 Compute Budget and Settings
Uniform wall-clock budget per issue (e.g., 60–120 minutes), identical tool access, and fixed seeds are used across compared systems where applicable. Summary statistics for the OSS study are reported in Table 7.
6.4 Repository-level OSS study
In addition to public benchmarks, an exploratory repository-level study is conducted on 10 open-source Python repositories spanning low- to high-complexity maintenance settings. A total of 100 issues (10 per repository) represents a spectrum of complexity, from low-dependency utilities (jsonpickle) to high-complexity image processing libraries involving C-extensions (Pillow). The selection criteria required repositories to have active maintenance, >1k>1k stars, and a history of well-documented issue resolution to serve as the ground truth. ”Complexity” is qualitatively assessed based on dependency depth and code coupling (see Table 3). The purpose of this study is ecological validity; it is not intended to replace standardized benchmarks such as SWE-bench.
Each system is evaluated under the same wall-clock and tool budget. A task in each trajectory is counted as resolved only if the generated patch satisfies the issue requirements under manual inspection and all repository tests pass. Because the issue set is curated for resolvability and the context is reconstructed from the final PR, these results should be interpreted as a controlled exploratory comparison rather than as a definitive leaderboard.
Table 3: Open-source repositories used in the OSS study (snapshot: 2025-12-05).
@@ 6.4 Repository-level OSS study
    Reported %Resolved on SWE-bench at time of writing [17] (original data in [21]).
@@
    For current SWE-bench leader-boards and variants, see the website [17, 14].
@@ 7.1 Components
The ablation study isolates (i) RL control (AIR w/o RL), and (ii) DTG (AIR w/o Graph). AIR w/o RL removes only the learned control policy while keeping the same DTG-based localization, editing, and validation pipeline. AIR w/o Graph removes the DTG representation and runs the same repair loop in a file-first workspace. The comparison is intended to separate the effect of repository representation from the effect of navigation/control.
@@ 8.1 Ablations: What matters and why
Table 5 isolates the contributions of the two main design choices. Removing RL control reduces %Resolved from 87.1 to 82.5. Although MTTR becomes shorter, this is not a gain in efficiency: the agent terminates earlier and explores fewer candidate repair trajectories. Removing DTG causes a much larger drop, to 56.2, indicating that repository representation is the dominant factor for multi-file fault localization. In other words, RL improves search control, but DTG provides the causal structure that makes the search tractable in the first place.
Overall, RL and Graph are the largest drivers of the resolution rate.
@@ 8.2 Comparison with commercial tools
Table 6 and Table 7 summarize model performance across 10 OSS projects (100 total issues). This study is less rigorous than SWE-bench for two reasons. First, the sample size is smaller. Second, 10 issues per repository were pre-selected for likely resolvability, which introduces selection bias toward coding issues that can be addressed through standard debugging and repair practice.
For each case, the evaluation context was reconstructed from the final pull request that resolved the issue.
A score of +1 is assigned when the issue requirements are satisfied and all repository unit tests pass. No penalty is assigned for time, cost, or failed attempts. Figure 3 shows the per-project comparison; the full results are reported in Table 6, with summary statistics in Table 7.
@@
Figure 3: Issue-resolution results on top 5 OSS repositories (10 issues per repository).
Table 6: Issue resolution on ten OSS GitHub repos (10 issues each; same wall-clock + tool budget). Numbers indicate how many issues were resolved.
@@
Table 7: Performance statistics per model across 10 projects (100 total issues).
@@
  • AIR achieved the highest score among the compared systems, solving 70% of all issues. It was the only model to solve at least 4 issues in every project.
@@ 8.2.1 Performance Analysis and the ”Semantic Gap”
The results in Figure 3 and Table 6 reveal a strong correlation between the explicitness of the data flow and the agent’s success rate. AIR achieves a 70% resolution rate overall, showing particular strength in ”loose-coupling” architectures like requests and urllib3 (8/10 and 9/10 respectively). In these repositories, logic errors are often distributed across multiple abstraction layers - a scenario where the DTG’s ability to trace data lineage allows the agent to ”hop” over irrelevant middleware, a capability lacking in the file-based baselines.
@@ 8.2.2 Failure Case Study: The Pillow Repository
A notable deviation is observed in the Pillow repository, where commercial models scored 0/10 and AIR achieved only 4/10. This performance drop stems from the ”Language Boundary” problem. Pillow relies heavily on C-extensions for image processing. The current DTG implementation parses Python ASTs but treats C-bindings as opaque edges. Consequently, when data flows into a compiled extension, the lineage is broken, blinding the agent. This limitation highlights the necessity for future work in cross-language graph construction to support hybrid-runtime environments.
9 Limitations
The present results should be interpreted with three limitations. First, the current DTG builder primarily targets Python and treats cross-language boundaries (e.g., C/C++ extensions) conservatively, which reduces observability in mixed-language repositories. Second, the OSS study uses a curated issue set and reconstructed context, so it is less rigorous than standardized benchmarks and may overestimate deployment-time performance. Third, AIR is evaluated empirically; the current version does not provide formal correctness guarantees beyond repository tests and static/dynamic validation signals. Consequently, AIR should be viewed as a practical empirical repair system rather than a formally verified maintenance framework.
10 Conclusion and Future Work
This paper has argued that the dominant file-first, control-centric view of repositories is misaligned with both human debugging practice and the causal structure of many real bugs. By re-centering our representation around data states and transformations, the Data-First Transformation Graph (DTG) offers a compact and causally meaningful substrate for LLM-based repair agents.
The empirical results on SWE-bench Verified, where AIR achieved an 87.1% resolution rate, validate that shifting from text-based retrieval to graph-based navigation substantially reduces the search complexity of logic repair. However, our analysis of failure cases, particularly within the Pillow repository, highlights critical avenues for future development. We outline three primary directions for the evolution of AIR:
1. Breaking the Language Boundary (Polyglot DTGs): The experiments revealed that the current DTG construction, while effective for pure Python, struggles with hybrid runtimes (e.g., Python wrappers around C/C++ extensions). Future iterations must implement cross-language graph builders that can trace the data lineage through Foreign Function Interfaces (FFI). This involves unifying AST parsers across languages into a shared intermediate representation, allowing the agent to ”tunnel” through compiled extensions rather than treating them as opaque edges.
2. Dynamic Graph Refinement via RL: Currently, the DTG is static, constructed prior to the agent’s traversal. This paper proposes a dynamic architecture where the agent can modify the graph structure at runtime—collapsing well-tested subgraphs to save context or expanding ambiguous nodes for deeper inspection. By integrating this into the Reinforcement Learning policy, the agent can learn to optimize its own environment representation, effectively performing ”attention management” over the repository topology.
@@ 10 Conclusion and Future Work
Overall, AIR suggests that data-lineage-aware navigation can improve repository -scale repair when compared to file-first interaction alone. The main contribution of this work is therefore not a claim of universal autonomy but a concrete data-first representation and agent design that improves causal localization under repository-scale context constraints. The results are encouraging, while the identified limitations define the work still required for a larger deployment.
10.0.1 Acknowledgments.
The author thanks Modal and lightning.ai for providing compute resources and free cloud credits for our experiments.
Appendix 1
@@ 27 ConsoleMessage(
Listing 1: A code snippet from poetry project - some non-essential lines were cut for brevity (poetry/ src/poetry/installation/chooser.py:125:197)
The graph builder can extract the following raw constructs from this snippet of code (Table 8) and build the subgraph depicted in Fig. 4. The standard Python ast library is used for parsing Python files, Neo4j for storing nodes and edges, and ChromaDB as a vector-storage tier for semantic search over relevant nodes and edges.
@@
Table 8: Summary of extracted code constructs. Verbose metadata such as file name, line number, and variable type is omitted for brevity. Function arguments, return values, and local variables become data nodes. The function name forms the basis for the corresponding transformation edge. The resulting structures are stored in Neo4j, while code embeddings are stored in ChromaDB.
@@
Figure 4: The reconstructed sample of Data Transformation Graph given the raw data to build Nodes and Edges from Table 8.
Appendix 2
Table 9 presents a comparison of the market tier, interaction modality, and primary capabilities of current AI coding tools. AIR is positioned differently in its focus on asynchronous, autonomous maintenance.
Table 9: Representative AI-assisted software tools (Q4 2025). The comparison distinguishes market tier, interaction modality, primary capability, and operational scope relevant to software maintenance automation.
System Tier Modality Primary capability Scope / distinguishing characteristic
@@
GitHub Copilot Professional IDE/editor plugin Code completion, chat, refactoring Strong ecosystem integration; optimized for interactive developer assistance.
@@
Amazon CodeWhisperer Professional IDE plugin Code completion, security hints Security-oriented suggestions integrated with cloud development workflows.
@@
Tabnine Professional IDE plugin Local/cloud completion Emphasis on privacy-preserving deployment and enterprise control.
@@
Cursor Professional AI-first IDE Conversational editing, repository search Editor centered on agentic edits with strong repository-context support.
@@
Claude Code Professional Chat + tools Conversational code reasoning Long-context code reasoning with strong natural-language interaction.
@@
OpenAI Codex (legacy) Professional API/chat (retired) Code generation foundation model Historical foundation system that influenced later tool development.
Sourcegraph Professional IDE/plugin platform Repository reasoning, code editing Broad repository navigation and assistant functionality across large codebases.
B44 (builder) Basic Web builder Prompt-based app scaffolding Rapid MVP creation; limited support for repository-scale maintenance.
@@
Lovable Basic Web builder No/low-code generation Accessible rapid prototyping; limited support for multi-file change management.
Replit (Ghostwriter) Basic Web IDE In-editor assistance Lightweight assistance oriented toward education and small-team workflows.
AlphaCode [11] Specialized Research system Competition-style code generation Strong performance on competitive programming tasks; not designed for repository maintenance.
@@
AlphaEvolve [13] Specialized Research system Algorithm discovery/optimization Effective in narrow optimization settings; not a general software maintenance system.
@@
Diffblue Cover Specialized JVM toolchain Unit test synthesis Industrial-grade test generation; complementary to, rather than a substitute for, issue repair.
@@
Getafix/Infer lineage [5] Specialized Internal/OSS tools Learned fix patterns, static analysis Pattern-based repair pipeline requiring suitable bug/fix corpora and analysis infrastructure.
@@
Factory.ai Specialized Cloud integration Agent-native software development Enterprise-oriented maintenance workflow support.